Rocephin (Ceftriaxone) 975 is ordered as a one time dose. 2gm of Rocephin reconstituted with 19.2 mL of dilutent is equal to 150 mg/mL. How many mL should be withdrawn from the vial? Round to the nearest tenth
Answer #1

Ordered Rocephin (Ceftriaxone)= 975 is ordered as a one time dose.

Instructions are to add 19.2 ml of diluent to yield 150 mg/ml

Means 150 mg in 1 ml

then 1 mg will contain =1/150 ml

for 975 mg required ml to be withdrawn= (1/150) X975= 6.5 ml

Thus 6.5 ml should be withdrawn from the vial after reconstituting the drug as isstructed.